# Psycholinguistic Analyses in Software Engineering Text: A Systematic Literature Review

### v1

@@ 5.5.1 Concerns with LIWC’s inherent limitations as a lexicon-based tool
A different limitation of all psycholinguistic tools, is that the author of any text might intentionally attempt to deceive the tool. Any individual may decide to purposefully use a specific type of language to be perceived in a particular way. However, the study bringing up this concern dismisses the potential concern by explaining that their collected data from GitHub contains communications among users who were not aware of the fact that their messages will be psycholingustically analyzed [S7].
@@ 5.5.2 Concerns with LIWC’s adoption in SE
Additionally, it can be argued that, so far the assessment of LIWC’s direct interpretability for psychological concepts and human factors has been mostly limited to general text. Therefore, to genuinely assess the tool’s reliability for domain-specific text, it is imperative to replicate studies exploring the correlations between LIWC’s output and psychological concepts using software engineering text such as, developers’ mailing lists and code reviews. Such an approach could help mitigate the performance loss in tasks such as personality detection, that involve understanding domain-specific communication nuances [S19].
@@
[backgroundcolor=lightgray!20,topline=false,leftline=false,rightline=false,bottomline=false] RQ5 Summary: We identify two types of concerns with the use of LIWC in SE: (a) inherent limitations due to its lexicon-based nature (e.g., lack of negation handling, limited language support), and (b) concerns with domain adoption (e.g., absence and potential of misrepresentation of SE-specific words).
@@ 6 Discussion
In this section, we explore the possibilities of improving and utilizing LIWC in novel ways for software engineering research.
@@ 6.1 Software engineering specific dictionaries
As detailed in Section 5.5.1 certain concerns associated with LIWC are inherent in any lexicon-based approach while others are the result of adopting the tool in a new field. Among the latter concerns, the issue often brought up in the literature is the domain specificity of software engineering text and LIWC’s limitations in capture the text’s context. The answer to this problem may lie in the creation of software-specific LIWC dictionaries which can prove to be a huge step forward toward a better understanding of developers and their unique linguistic patterns. While creating any LIWC dictionary demands a notable investment of time and effort, once created, it can present new opportunities for those studying developers’ communications. Such dictionaries can be specific to particular settings in order to capture the language patterns in certain situations. For example, researchers interested in investigating the levels of stress and anxiety among developers could develop a custom dictionary that reflects the language use of developers in high-pressure situations. Overall, LIWC possesses the potential to significantly improve our understanding of the text written by developers and therefore, addressing some of LIWC’s limitations can prove to be invaluable.
@@ 6.2 Leveraging more LIWC categories for direct analysis of the text
Another way to improve the usability of LIWC, is to take advantage of all its categories for direct analytical purposes. LIWC presents an expansive set of word categories. However, commonly, studies select specific categories, tailored to the needs of their particular analytical objectives, for close examination. As evident in Table LABEL:tab:all_categories, our findings point to the diversity of LIWC categories that have been utilized to serve different research purposes in software engineering studies. Yet, even though many software engineering studies have leveraged several of LIWC’s capabilities, we still believe that the program holds untapped potential for conducting further text analysis through direct interpretation of its categories. Consider the category of ‘assent’, known to correlate with agreement and passivity [22]. Understanding assent can potentially help us in studying team conflicts, or analyzing discussions involving critical decision-making, such as tool adoption. Despite its potential relevance, this category has yet to be employed in order to assess agreement levels within developer conversations.
@@ 6.3 LIWC and Large Language Models
To overcome the interpretability challenges posed by black-box models, one promising approach involves integrating LIWC to enhance their ability to analyze text. As highlighted in our response to RQ1, LIWC categories have already been utilized as input in various machine learning models, often in combination with other features such as metadata from sources like Stack Overflow posts or GitHub repositories. By leveraging LIWC’s psychologically grounded insights, deep learning models could be enriched with valuable context that improves their ability to capture nuanced human emotions and cognitive states. Given the challenges in interpretability that many deep learning models face [82], this hybrid approach offers a promising direction for future research and application.
Additionally, there is considerable room for analyzing LLM-generated content with the help of LIWC. Beyond evaluating the content of ChatGPT-generated responses, Kabir et al. analyzed the linguistic characteristics of such content generated in response to Stack Overflow questions and compared them to those written by humans [S43]. Their findings revealed that LLM-generated content tends to be more formal, analytic, and positive in language. This kind of analysis can be extended to explore the linguistic features of different LLMs in various scenarios and conversations in software contexts. For instance, it would be insightful to determine what type of language is most effective in prompting an LLM when trying to solve programming issues or address GitHub issues in a repository. Furthermore, it would be valuable to study how LLMs navigate conversations based on specific linguistic cues, whether they adjust their style in certain contexts, and how these adjustments impact the usefulness of their responses.
@@
6.4 New Research Opportunities Leveraging LIWC
@@ 7 Threats to Validity
Construct Validity:
Papers retrieved from online libraries were manually filtered to select those pertinent to this study. While this method can be susceptible to human error, we have implemented a meticulous two-step process to mitigate this risk. In addition to assessing the title and abstract, we thoroughly review the entire paper before determining its inclusion in our study. This approach ensures the precision and relevance of our paper selection.
Internal Validity:
The internal operations of the online libraries we utilized in our study are undisclosed. Their methods of indexing studies and responding to search queries are not publicly available. Therefore it is possible, that this search process has potentially resulted in the omission of relevant studies. However, to mitigate this threat, we have collected relevant papers from six prominent and widely recognized online digital libraries, i.e., IEEE Xplore Digital Library, ACM Digital Library, Elsevier Science Direct, Springer Link Online Library, Wiley Online Library, and Scopus.
External Validity: In order to narrow down the survey to relevant papers, we implemented stringent inclusion and exclusion criteria. This led to our focus on papers written in English and accessible through one of our six chosen databases. However, this approach may have missed some papers, potentially affecting the study’s generalizability. To mitigate this risk, we designed our search string to be as inclusive as possible, deliberately introducing more noise into our initial results. We later performed a meticulous manual analysis to eliminate irrelevant entries. This strategy aimed to prioritize recall over precision, ensuring the inclusion of as many relevant papers as possible. Consequently, we invested significant time in the manual inspection process, resulting in the removal of over a thousand irrelevant papers.
@@ 8 Conclusion
In addressing our first research question, we underlined the diverse purposes for which LIWC has been used within the software engineering domain. Our findings reveal an almost even split between studies that directly interpret LIWC’s various dimensions and those that employ it in an indirect manner, often combined with other data or as input for machine learning models. We also highlight how, throughout the years, the community has demonstrated many unique methods of leveraging LIWC to address development-related issues and enhance the software engineering process. Furthermore, we present the types of data and data sources that were utilized in our selected studies, and leverage a taxonomy of software engineering activities adopted from previous research, to identify the activities that were most frequently studied using LIWC (e.g., Team Management and Knowledge Management). Additionally, by adopting Behavioral Software Engineering (BSE) concepts, we carefully annotated each paper based on the concepts it examined, offering a precise overview of the distribution of these studied concepts. This step allowed us to identify the concepts that were frequently studiedusing LIWC, as well as the ones that have received less attention or remained unexplored. Next, to better understand the potential and limitation of LIWC in analyzing domain-specific language used by software engineering practitioners, we investigated the evaluation methods used by researchers. Surprisingly, we discovered that a substantial portion of the papers neither directly evaluated LIWC’s findings nor raised any concerns regarding its usage. Nonetheless, a noteworthy number of papers do subject LIWC’s findings to different types of evaluations.
@@ Appendix A Selected Studies
Below you can find a list of the studies used in this systematic literature review:
@@
[37] Silva, L. et al. (2022) ‘Emotional Dashboard: a Non-Intrusive Approach to Monitor Software Developers’ Emotions and Personality Traits’, in 2022 IEEE 22nd International Conference on Software Quality, Reliability and Security (QRS). IEEE, pp. 366–375.

### v2

@@ 5.5.1 Concerns with LIWC’s inherent limitations as a lexicon-based tool
A potential concern with psycholinguistic tools is that authors might intentionally manipulate their language to influence the analysis. However, this is largely irrelevant in SE contexts, where communications, such as those on GitHub, are organic and not written with the expectation of being psycholinguistically analyzed [S7]. As such, we do not consider this a meaningful limitation for our review.
@@ 5.5.2 Concerns with LIWC’s adoption in SE
Additionally, it can be argued that, so far the assessment of LIWC’s direct interpretability for psychological concepts and human factors has been mostly limited to general text. Therefore, to genuinely assess the tool’s reliability for domain-specific text, it is imperative to replicate studies exploring the correlations between LIWC’s output and psychological concepts using software engineering text such as, developers’ mailing lists and code reviews. Such an approach could help mitigate the performance loss in tasks, such as personality detection, that involve understanding domain-specific communication nuances [S19].
Although our review primarily focused on concerns focused on LIWC’s use in software engineering, some studies have also noted broader ethical considerations when employing psychometric tools in this domain. For example, S19 acknowledges concerns regarding the use of personality detection techniques in SE research. While such concerns are not specific to LIWC itself, the use of automated psycholinguistic tools can raise important questions around privacy, consent, and the interpretation of psychological inferences. We acknowledge the importance of these issues and note that researchers should exercise care when applying such tools to sensitive or personal aspects of developer behavior.
@@
[backgroundcolor=lightgray!20,topline=false,leftline=false,rightline=false,bottomline=false] RQ5 Summary: We identify two types of concerns with the use of LIWC in SE: (a) inherent limitations due to its lexicon-based nature (e.g., lack of negation handling, limited language support), and (b) concerns with domain adoption (e.g., absence and potential misrepresentation of SE-specific words).
@@ 6 Discussion
Overall, our review indicates that LIWC has been successfully applied across a broad spectrum of SE tasks involving human factors, such as analyzing team dynamics, detecting emotions, and investigating various BSE concepts (e.g., leadership and personality). Studies have also demonstrated how some psycholinguistic frameworks from other domains that leverage LIWC can be adopted in SE studies, enriching SE research. However, several weaknesses were identified as well. Most notably, few papers rigorously evaluated the validity of LIWC’s outputs for their specific SE contexts; many appeared to assume that the tool’s general reliability would naturally extend to SE text. Furthermore, our findings show that no effort has been made to develop SE-specific dictionaries, which could address domain-related limitations and improve LIWC’s interpretability for SE-specific text. As a result, more validation efforts are needed to ensure that LIWC’s psychological categories can accurately assess SE text. In the remainder of this section, we explore some possibilities for improving and utilizing LIWC in novel ways for software engineering research.
@@ 6.1 Software engineering specific dictionaries
As detailed in Section 5.5.1 certain concerns associated with LIWC are inherent in any lexicon-based approach while others are the result of adopting the tool in a new field. Among the latter concerns, the issue often brought up in the literature is the domain specificity of software engineering text and LIWC’s limitations in capturing the text’s context. The answer to this problem may lie in the creation of software-specific LIWC dictionaries which can prove to be a huge step forward toward a better understanding of developers and their unique linguistic patterns. While creating any LIWC dictionary demands a notable investment of time and effort, once created, it can present new opportunities for those studying developers’ communications. Such dictionaries can be specific to particular settings in order to capture the language patterns in certain situations. For example, researchers interested in investigating the levels of stress and anxiety among developers could develop a custom dictionary that reflects the language use of developers in high-pressure situations. Furthermore, custom dictionaries could be defined with specific software engineering constructs, such as Team Management or Quality Assurance, in mind to more accurately capture the language patterns and psychological markers relevant to each activity. For example, dictionaries designed around team coordination could emphasize constructs like leadership, trust, or conflict resolution, enhancing the interpretability of LIWC’s output in those contexts. Overall, LIWC possesses the potential to significantly improve our understanding of the text written by developers and therefore, addressing some of LIWC’s limitations can prove to be invaluable.
@@ 6.2 Leveraging more LIWC categories for direct analysis of the text
Another way to improve the usability of LIWC is to take advantage of all its categories for direct analytical purposes. LIWC presents an expansive set of word categories. However, commonly, studies select specific categories, tailored to the needs of their particular analytical objectives, for close examination. As evident in Table LABEL:tab:all_categories, our findings point to the diversity of LIWC categories that have been utilized to serve different research purposes in software engineering studies. Yet, even though many software engineering studies have leveraged several of LIWC’s capabilities, we still believe that the program holds untapped potential for conducting further text analysis through direct interpretation of its categories. Consider the category of “assent”, known to correlate with agreement and passivity [22]. Understanding assent can potentially help us in studying team conflicts, or analyzing discussions involving critical decision-making, such as tool adoption. Despite its potential relevance, this category has yet to be employed in order to assess agreement levels within developer conversations.
@@ 6.3 LIWC and Large Language Models
To overcome the interpretability challenges posed by black-box models, one promising approach involves integrating LIWC to enhance their ability to analyze text. As highlighted in our response to RQ1, LIWC categories have already been utilized as input in various machine learning models, often in combination with other features such as metadata from sources like Stack Overflow posts or GitHub repositories. By leveraging LIWC’s psychologically grounded insights, deep learning models could be enriched with valuable context that improves their ability to capture nuanced human emotions and cognitive states.
To overcome the interpretability challenges posed by black-box models, one promising approach involves integrating LIWC to enhance their ability to analyze text. As highlighted in our response to RQ1, LIWC categories have already been utilized as input in various machine learning models, often in combination with other features such as metadata from sources like Stack Overflow posts or GitHub repositories. LIWC features could even be incorporated into ensemble classifiers to improve performance across SE-related prediction tasks. By leveraging LIWC’s psychologically grounded insights, deep learning models could also be enriched with valuable context that improves their ability to capture nuanced human emotions and cognitive states. Given the challenges in interpretability that many deep learning models face [90], this hybrid approach offers a promising direction for future research and application.
Additionally, there is considerable room for analyzing LLM-generated content with the help of LIWC. Beyond evaluating the content of ChatGPT-generated responses, Kabir et al. analyzed the linguistic characteristics of such content generated in response to Stack Overflow questions and compared them to those written by humans [S43]. Their findings revealed that LLM-generated content tends to be more formal, analytic, and positive in language. These distinctions also suggest the potential for using LIWC to help differentiate between human- and LLM-authored text in software engineering artifacts. This kind of analysis can be extended to explore the linguistic features of different LLMs in various scenarios and conversations in software contexts. For instance, it would be insightful to determine what type of language is most effective in prompting an LLM when trying to perform programming tasks or address GitHub issues in a repository. Furthermore, it would be valuable to study how LLMs navigate conversations based on specific linguistic cues, whether they adjust their style in certain contexts, and how these adjustments impact the usefulness of their responses.
@@
7 New Research Opportunities Leveraging LIWC
@@
Finally, beyond the immediate applications listed above, future studies can leverage LIWC in multiple ways to further strengthen behavioral software engineering research. For instance, developing custom SE-specific dictionaries around targeted psychological constructs or BSE concepts (see Section 6.1) could provide more precise insights into team dynamics or leadership behaviors. Researchers might also explore underutilized LIWC categories (as discussed in Section 6.2) to capture new dimensions of developer communication that have not yet been investigated. Likewise, combining LIWC with deep learning or language models (Section 6.3) may yield both improved classification performance and more interpretable results in tasks such as developer sentiment analysis or automated code review. By expanding these efforts, the community can continue pushing the boundaries of LIWC’s utility, creating a deeper, more contextualized understanding of human factors in software engineering.
@@ 8 Threats to Validity
Construct Validity: Papers retrieved from online libraries were manually filtered to select those pertinent to this study. While this method is susceptible to human error, we adopted a two-step process: one author conducted the primary selection, while another independently verified the final list. Although this adds a layer of validation, it does not fully replicate having two authors review every paper, and some risk of missing relevant studies remains. In terms of our search query, a potential threat to validity is the assumption that the naming conventions for LIWC are consistent across all relevant literature, which may overlook studies using alternate terminology. We mitigated this threat by explicitly checking for alternate naming conventions during our literature review but found none in the relevant studies across software engineering, psychology, and communication domains. In addition to assessing the title and abstract, we thoroughly review the entire paper before determining its inclusion in our study. This approach ensures the precision and relevance of our paper selection. To further mitigate bias, two authors were involved in the data extraction and coding phases of the study. We followed an iterative discussion and coding procedure until we reached substantial agreement for each category (above 0.6).
Internal Validity: The internal operations of the online libraries we utilized in our study are undisclosed. Their methods of indexing studies and responding to search queries are not publicly available. Therefore it is possible, that this search process has potentially resulted in the omission of relevant studies. However, to mitigate this threat, we have collected relevant papers from six prominent and widely recognized online digital libraries, i.e., IEEE Xplore Digital Library, ACM Digital Library, Elsevier Science Direct, Springer Link Online Library, Wiley Online Library, and Scopus.
External Validity: In order to narrow down the survey to relevant papers, we implemented stringent inclusion and exclusion criteria. This led to our focus on papers written in English and accessible through one of our six chosen databases. However, this approach may have missed some papers, potentially affecting the study’s generalizability. To mitigate this risk, we designed our search string to be as inclusive as possible, deliberately introducing more noise into our initial results. We later performed a meticulous manual analysis to eliminate irrelevant entries. This strategy aimed to ensure the inclusion of as many relevant papers as possible. Consequently, we invested significant time in the manual inspection process, resulting in the removal of over a thousand irrelevant papers.
@@ 9 Conclusion
In addressing our first research question, we underlined the diverse purposes for which LIWC has been used within the software engineering domain. Our findings reveal an almost even split between studies that directly interpret LIWC’s various dimensions and those that employ it in an indirect manner, often combined with other data or as input for machine learning models. We also highlight how, throughout the years, the community has demonstrated many unique methods of leveraging LIWC to address development-related issues and enhance the software engineering process. Furthermore, we present the types of data and data sources that were utilized in our selected studies, and leverage a taxonomy of software engineering activities adopted from previous research, to identify the activities that were most frequently studied using LIWC (e.g., Team Management and Knowledge Management). Additionally, by adopting Behavioral Software Engineering (BSE) concepts, we carefully annotated each paper based on the concepts it examined, offering a precise overview of the distribution of these studied concepts. This step allowed us to identify the concepts that were frequently studiedusing LIWC, as well as the ones that have received less attention or remained unexplored. Next, to better understand the potential and limitation of LIWC in analyzing domain-specific language used by software engineering practitioners, we investigated the evaluation methods used by researchers. Surprisingly, we discovered that a substantial portion of the papers neither directly evaluated LIWC’s findings nor raised any concerns regarding its usage. Nonetheless, a noteworthy number of papers do subject LIWC’s findings to different types of evaluations. We also note that ethical considerations, such as privacy and the potential for misinterpretation of psychometric inferences, are rarely addressed.
@@ Appendix A Selected Studies
Below you can find a list of the studies used in this SLR:
@@
[S37] Licorish, S. A. and MacDonell, S. G. (2021) ‘Relating IS Developers’ Attitudes to Engagement’, in Proceedings of the 25th Australasian Conference on Information Systems (ACIS’14). 1–10.
@@
Appendix B PRISMA 2020 Items
Table 10: PRISMA 2020 Items with Reporting Locations
Checklist Item Location Reported
Title: Identify the report as a systematic review. Title
Abstract: See the PRISMA 2020 for Abstracts checklist. Abstract
Introduction – Rationale: Describe the rationale for the review in the context of existing knowledge. Section 1
Introduction – Objectives: Provide an explicit statement of the objectives or questions the review addresses. Section 1
Methods – Eligibility criteria: Specify inclusion and exclusion criteria and how studies were grouped. Section 4.2
Methods – Information sources: Specify all sources searched and date of last search. Sections 4.1 & 4.2
Methods – Search strategy: Present full search strategies for all sources. Section 4
Methods – Selection process: Describe methods for deciding study inclusion and number of reviewers involved. Section 4
Methods – Data collection process: Specify methods used to collect data. Section 4.3
Methods – Data items (Outcomes): List and define all outcomes for which data were sought. Section 4
Methods – Data items (Other variables): List and define all other variables for which data were sought. N/A
Methods – Study risk of bias assessment: Methods used to assess risk of bias. Section 4
Methods – Effect measures: Effect measures used for each outcome. N/A
Methods – Synthesis methods: Describe methods used to synthesize results and explore heterogeneity. Section 4
Methods – Reporting bias assessment: Describe methods used to assess risk of bias due to missing results. N/A
Results – Study selection: Describe results of search and selection. Section 5
Results – Excluded studies: Cite studies that may appear to meet inclusion criteria but were excluded. Replication Package
Results – Study characteristics: Cite each included study and present characteristics. Appendix A and Section 5
Results – Results of individual studies: Present summary statistics and effect estimates. N/A
Results – Results of syntheses: Present results of all statistical syntheses, including heterogeneity. N/A
Discussion – Summary of evidence: General interpretation of the results in the context of other evidence. Section 6
Discussion – Limitations of evidence: Discuss limitations of included studies. Section 6
Discussion – Limitations of review process: Discuss limitations of the review process. 8
Discussion – Implications: Discuss implications for practice, policy, and future research. Section 6
Other – Availability of data: Report availability of data, code, and materials. Section 4.3